# Leakage and Interpretability in Concept-Based Models

### v1

@@ Abstract
Concept Bottleneck Models aim to improve interpretability by predicting high-level intermediate concepts, representing a promising approach for deployment in high-risk scenarios. However, they are known to suffer from information leakage, whereby models exploit unintended information encoded within the learned concepts. We introduce an information-theoretic framework to rigorously characterise and quantify leakage, and define two complementary measures: the concepts-task leakage (CTL) and interconcept leakage (ICL) scores. We show that these measures are strongly predictive of model behaviour under interventions and outperform existing alternatives in robustness and reliability. Using this framework, we identify the primary causes of leakage and provide strong evidence that Concept Embedding Models exhibit substantial leakage regardless of the hyperparameters choice. Finally, we propose practical guidelines for designing concept-based models to reduce leakage and ensure interpretability.
@@ 1 Introduction
Explainability and transparency are essential aspects of model design, especially in high-risk scenarios such as biomedical applications. Concept Bottleneck Models (CBMs) (Kumar et al., 2009; Lampert et al., 2009; Koh et al., 2020) stand out in the landscape of interpretable models as they learn a set of high-level intermediate concepts associated with an outcome, in a manner which facilitates human oversight (Banerji et al., 2025). CBMs first predict the level of activation of each concept for a given input data point, then use such activations to predict a task. Their structure makes them self-explainable, and allows for direct human supervision and intervention, as opposed to post hoc interpretability approaches.
@@
A trained concept-based model, however, may be significantly less interpretable than expected, despite presenting the illusion of being so. A key reason for this is the phenomenon of information leakage (Kazhdan et al., 2021; Margeloiu et al., 2021; Mahinpei et al., 2021; Havasi et al., 2022). Leakage occurs when a poorly designed concept-based model uses additional information (such as interconcept relationships) to attain a higher task accuracy than a well-designed (i.e. transparent) concept-based model, at the cost of undermining concept interpretability. Storing such additional information into the learnt concepts is beneficial for the model with respect to loss function minimisation, as it typically allows the final head to leverage this information, attaining higher task performance while preserving high accuracy in concept predictions. This ultimately results in learnt concepts not being aligned with the ground-truth concepts, thus hindering the actual interpretability of the model. For example, in the clinical setting of cancer diagnosis, the grade of a tumour (how aggressive it looks under a microscope) is correlated to tumour stage (how much the cancer has spread through the body) and leakage may occur between learned concept representations of grade and stage. This means that a concept representation of grade might also contain information about tumour stage. However, these concepts have independent clinical definitions and their appropriate assessment is essential to guide clinical management. A concept-based model which displays leakage may therefore be unsafe for critical decision-making applications, such as clinical scenarios.
@@
We then employ these measures to identify and assess the primary causes of leakage in CBMs, including over-expressive concept representations, an incomplete set of annotated concepts, a misspecified final head, and insufficient concept supervision. Building on this novel information-theoretic framework, we provide strong evidence that CEMs and closely related models exhibit dramatic leakage regardless of the choice of hyperparameters, thereby casting serious doubts on their interpretability. Finally, we propose a set of guidelines for the design of concept-based models based on preventing leakage, and we argue that evaluating leakage is a crucial step of model development to ensure interpretability.
Note: During the final stages of this work, Makonnen et al. (2025) appeared, presenting minor overlap with the content discussed here.
@@ 2 Concept bottleneck models
Consider a dataset $\left\{x^{(n)},c^{(n)},y^{(n)}\right\}_{n=1}^{N}$ with $N$ observations, where $x^{(n)}\in\mathbb{R}^{d}$ is the input, $c^{(n)}\in\{0,1\}^{k}$ represents a set of $k$ annotated concepts, and $y^{(n)}\in\{1,\dots,\ell\}$ is the target label. A CBM (Figure 1) is the composition of a concept encoder $g^{*}$ mapping the input to the predicted concept activations $\hat{c}=g^{*}(x)$, and a classifier head $f^{*}$ which maps the concepts to the predicted label $\hat{y}=f^{*}(\hat{c})$. This framework easily extends to regression tasks $y$ and to categorical or continuous concepts.
The presence of a concept bottleneck displaying the predicted concept activations for each input allows a human agent to directly supervise the model reasoning without resorting to post-hoc approaches to explainability such as CAM (Zhou et al., 2016), GradCAM (Selvaraju et al., 2017), LIME (Ribeiro et al., 2016) or SHAP (Lundberg and Lee, 2017). Additionally, this setup enables human interventions: if one or more predicted concepts are identified as incorrect for a given input $x^{(n)}$, a human agent can fix their values according to their expertise, resulting in an improved task prediction.
To achieve full interpretability, the final head architecture should represent a functional dependence that is interpretable. This allows for a higher level of supervision, e.g. in terms of which concepts are more relevant for the task prediction. A popular choice is linear classifiers.
@@
  • independent training, where the concept encoder and the final head are trained independently on ground-truth data,
@@
  • sequential training, where the concept encoder is trained on ground-truth data, while the final head is trained on the predicted concepts,
@@
  • joint training, where the model is trained end-to-end,
@@
    with $\lambda\geq 0$ hyperparameter indicating the importance of concept learning over task learning. $\lambda=0$ corresponds to a black box model.
@@
Common choices for concept encoding include binary probabilities, soft probabilities or logits. We refer to hard CBMs as independently trained models with binary concept encoding, while soft CBMs denote jointly trained models with soft probability-based concept activations. Additionally, we refer to jointly trained models using logit-based concept encoding as logit CBMs.
@@
More recently, CEMs (Figure 1) were proposed in Espinosa Zarlenga et al. (2022), where a pair of vectors $\left(\bm{\hat{c}^{+}}_{i},\bm{\hat{c}^{-}}_{i}\right)\in\mathbb{R}^{2d}$ is predicted for each input and concept $c_{i}$, $i=1\dots k$ (with $d$ hyperparameter indicating the embedding dimension). This pair of vectors is subsequently used to predict the activation $\hat{c}_{i}$ as a soft probability. The weighted vectors $\bm{\hat{c}^{w}}_{i}=\hat{c}_{i}\,\bm{\hat{c}^{+}}_{i}+(1-\hat{c}_{i})\,\bm{\hat{c}^{-}}_{i}$ are then constructed and concatenated before being fed into the final head to predict the label. In this setup $\bm{\hat{c}^{+}}_{i}$ and $\bm{\hat{c}^{-}}_{i}$ are meant to represent concept $i$ being active or inactive respectively. Learning a two-fold representation for each concept enables interventions, by modifying the concept activations $\hat{c}_{i}$ and thus selecting either $\bm{\hat{c}^{+}}_{i}$ or $\bm{\hat{c}^{-}}_{i}$ as an input for the final head.
@@
Training in CEMs is performed end-to-end; similar to joint CBMs, the loss function consists of reconstruction losses for both the concept activations and the class label, weighted by a parameter $\lambda\geq 0$ as in (3). To make such models more receptive to interventions (i.e. enhance the improvement in task performance after each intervention), they are exposed to interventions during training: each activation $\hat{c}_{i}$ is set to its ground-truth value $c_{i}$ with probability defined by a parameter $p_{int}\in(0,1)$.
@@
Reasoning in CEMs takes place at the level of the vectors $\bm{\hat{c}^{w}}_{i}$, as opposed to CBMs where it is based on concept activations $\hat{c}_{i}$. However, as high-dimensional objects trained only to be predictive of concepts $c_{i}\in\{0,1\}$, the vectors $\bm{\hat{c}^{w}}_{i}\in\mathbb{R}^{d}$ are excessively flexible representations, capable of encoding an arbitrary amount of information about the task label. We demonstrate in Section 8 that this is the fundamental reason why CEMs generally fail to be interpretable for any choice of their hyperparameters. While Espinosa Zarlenga et al. (2022) presents CEMs as interpretable models that resolve the accuracy-interpretability trade-off, they still adhere to this trade-off in favour of accuracy, effectively sacrificing interpretability. Although requiring a case-by-case evaluation, other CEM-based architectures, such as Espinosa Zarlenga et al. (2023b); Ridzuan et al. (2024); Gao et al. (2024); Hu et al. (2024); Srivastava et al. (2024), are likely to face similar interpretability challenges.
@@ 3 Definition of leakage
Figure 2: Distributions of predicted activations for concept 1 on the test set in CBMs with increasing leakage from left to right. Colours are based on the ground-truth values of concepts and task label.
@@
The explainability of CBMs and CEMs assumes that learnt concepts are aligned with ground-truth human-interpretable concepts. However, on general grounds that is not the case. During training, concept-based models often encode additional input information into the learned concepts to enhance task accuracy - information leakage occurs (Kazhdan et al., 2021; Margeloiu et al., 2021; Mahinpei et al., 2021; Havasi et al., 2022; Lockhart et al., 2022; Ragkousis and Parbhoo, 2024). The model is not relying solely on the value of concepts for its task prediction, but also on such leaked information, hindering interpretability. The reasoning of the model thus becomes effectively obscure to human agents, to a degree specified by the amount of leakage.
@@
In practice leakage manifests as additional structure arising in the learnt concept distributions, as showcased in Figure 2. In this example models must learn two binary concepts for a binary classification task (see Appendix A.1 for more details). We compare the distributions of activations for the first concept by a hard CBM and two soft CBMs, with either intermediate ($\lambda=1$) or low ($\lambda=0.01$) concept supervision. By definition no leakage can be present in hard CBMs, since concept learning is independent of task learning, and concepts are converted to binary values before being passed to the final head, preventing leaked information. Soft CBMs with lower concept supervision are instead more prone to leakage as accurate concept learning is not enforced during training (see Koh et al., 2020; Kazhdan et al., 2021, and our results in Section 7).
@@
Neglecting the peak close to 0.5 corresponding to observations the model is most uncertain about, the concept distribution in the hard CBM is very close to the binary ground-truth distribution for concept 1. In the soft CBM with $\lambda=1$ additional structure is learnt based on the value of the other concept and of the class label, while for $\lambda=0.01$ the concept itself is a very strong predictor of the value of both the class label and the other concept. In particular, the distribution is more similar to the ground-truth binary task distribution than to the concept distribution.
@@
We identify two main types of leakage that concept-based models resort to:
@@
1) Concepts-task leakage: additional information about the task is stored into the learnt concepts. For training to be successful, the annotated concepts must be predictive of the task label as measured by the ground-truth mutual information (MI) between each concept and the task label. If concepts-task leakage is present, learned concepts-task MI is higher than ground-truth concepts-task MI. This is usually the prominent type of leakage and the main cause of non-interpretability, as it directly allows the model to achieve a better task performance (see experiments in Section 6).
@@
2) Interconcept leakage: additional information about the other concepts is stored into each concept. In a given dataset, concepts are predictive of the value of other concepts as determined by their ground-truth pair-wise MI. This type of leakage manifests as learned concepts being more predictive of the value of the other learned concepts, resulting in a learned interconcept MI being higher than the ground-truth. This effect is usually a secondary factor of non-interpretability (see experiments in Section 6), but it may be beneficial for models as an internal error-correcting tool: interconcept leakage provides redundant pathways for the model to attain high task performance even when concept predictions are poor.
@@ 4 Existing measures of interpretability and their shortcomings
Previous works (Koh et al., 2020; Kazhdan et al., 2021; Margeloiu et al., 2021; Mahinpei et al., 2021; Havasi et al., 2022; Espinosa Zarlenga et al., 2023a) have proposed several measures of leakage, however none take into account the information-theoretic nature of leakage outlined in Section 3. As we highlight in this section, these existing measures consequently exhibit inherent limitations in sensitivity and robustness, resulting in a lack of applicability to real-world scenarios.
@@
Although performance metrics such as concept accuracy, F1 score and AUC are general quality indicators of concept learning, they are not sensitive to more subtle effects undermining interpretability such as leakage. In particular, two models may exhibit the same evaluation scores while considerably differing in terms of leakage. This is apparent from the example in Figure 2, where a classification threshold of 0.5 yields essentially the same concept accuracy and F1 score for the hard and $\lambda=1$ soft CBM, while the latter encodes a non-trivial amount of additional structure. Concept AUC is a finer measure of interpretability than accuracy and F1 score as it captures more information about the concept distribution, and in this simple example it is able to discriminate between those two models. However it is not generally sensitive enough to quantify leakage in more complex setups as we illustrate in Section 6.
@@
These scores were defined in Espinosa Zarlenga et al. (2023a). The Oracle Impurity Score (OIS) is obtained by training $k(k-1)$ additional neural networks (NNs) $\psi_{i,j}$ (typically 2-layer perceptrons of modest size) predicting the value of the ground-truth concept $j$ from the learnt representation of concept $i$. The $k\times k$ impurity matrix defined as ${\pi_{ij}(\hat{c},c)=AUC\left(\psi_{i,j}(\hat{c}_{i}),c_{j}\right)}$ is meant to estimate the predictivity of each learnt concept representation with respect to any other across the dataset. The OIS is the average difference between the predictivities of learnt and ground-truth concepts over pairs of concepts,
@@
where $||\cdot||_{F}$ indicates the Frobenius norm, and the normalisation ensures $0\leq\text{OIS}\leq 1$. The OIS is thus intended to provide an estimate of interconcept leakage, while being unable to account for concepts-task leakage, which is the prevalent effect as we illustrate in Section 6. Although demonstrated to be superior to other concept-quality metrics in Espinosa Zarlenga et al. (2023a), the OIS is subject to additional limitations that hinder its applicability. In particular, in Section 6 we prove that
@@
    its value is typically non-vanishing and relatively high for hard CBMs, indicating biases and a lack of sensitivity;
@@
Figure 3: Performance upon random intervention of the three models analysed in Figure 2.
The Niche Impurity Score (NIS) was defined to capture leakage manifesting as additional information about concept $j$ stored into a joint subset of the learnt representations $\{i_{1},\dots,i_{p}\}$ not containing concept $j$ (see Espinosa Zarlenga et al., 2023a, for details). Such a score is meant to go beyond pair-wise interconcept leakage and estimate the (generally subdominant) higher-order effects. As illustrated in Appendix B, the NIS is generally non-vanishing and very high for hard CBMs, and moreover it appears to be anticorrelated with intervention performance and leakage. As such, we deem it unsuitable to measure leakage and interpretability.
@@
Intervention performance can be measured at test time by substituting the ground-truth concept values to the predicted activations, with a given policy defining the order of the concepts to intervene on. The behaviour of the task accuracy after each intervention measures to what extent learnt and ground-truth concepts are aligned. A task accuracy that decreases is a coarse indication of leakage: the final head expects additional information that is not present in the ground-truth concept distribution. Figure 3 shows the intervention performance of the same three models as in Figure 2. Models with higher leakage have worse performance, while the hard CBM displays a monotonically increasing task accuracy.
As a metric of leakage, intervention performance has however several limitations.
@@
    Being based on evaluation scores, it does not address the information-theoretical nature of leakage, resulting in a limited resolution. It is unable to distinguish between interconcept and concepts-task leakage, and it does not directly quantify interpretability at the level of single concepts. Typically it is thus not sufficiently informative to guide model design and hyperparameter selection.
@@
    Its evaluation is computationally expensive, especially in real-world applications where models have a large number of concepts and are evaluated using state-of-the-art policies with a high overhead such as CooP (Chauhan et al., 2023).
@@
    In more complex architectures such as CEMs, it is not a measure of interpretability, see Appendix H.
Poor intervention performance is in particular a measure of leakage with perfect specificity (leakage has always occurred when performance is poor), but with poor sensitivity (when performance is good, leakage is not guaranteed to be absent). Although the leakage scores we propose in Section 5 address these issues, intervention performance generally remains a coarse but useful indicator of leakage in CBMs. Let us denote by $y_{acc}^{*(k)}$ the task accuracy after all the $k$ concepts have been intervened on in the model under evaluation, which coincides with the accuracy of the final head on the ground-truth concepts. As a benchmark of the proposed leakage scores we will use the intervention score
$S_{(int)}=y_{acc}^{(k)}-y_{acc}^{*(k)}\,,$ (5)
where $y_{acc}^{(k)}$ is the reference accuracy of the final head when separately trained on the ground-truth concepts. It represents the maximal task accuracy attainable by the final head and it measures the fundamental misspecification of the final head with respect to the ground-truth functional dependence $y=f(c)$. $S_{(int)}$ thus quantifies the further decrease in task accuracy caused by leakage on top of the possible final head misspecification. Its value is independent of the intervention policy adopted. Note that $y_{acc}^{(k)}$ coincides with $y_{acc}^{*(k)}$ for a successfully trained hard model, thus $S_{(int)}=0$ for hard models by construction.

### v2

@@ Abstract
Concept Bottleneck Models aim to improve interpretability by predicting high-level intermediate concepts, representing a promising approach for deployment in high-risk scenarios. However, they are known to suffer from information leakage, whereby models exploit unintended information encoded within the learned concepts. We introduce an information-theoretic framework to rigorously characterise and quantify leakage, and define two complementary measures: the concepts-task leakage (CTL) and interconcept leakage (ICL) scores. We show that these measures are strongly predictive of model behaviour under interventions and outperform existing alternatives in robustness and reliability. Using this framework, we identify the primary causes of leakage and provide strong evidence that Concept Embedding Models exhibit substantial leakage regardless of the hyperparameter choice. Finally, we propose a set of practical guidelines for designing concept-based models to reduce leakage and ensure interpretability.
Keywords: Concept Learning, Explainable AI (XAI), Interpretability metrics, Concept Bottleneck Models, Information leakage
@@ 1 Introduction
Explainability and transparency are essential aspects of model design, especially in high-risk scenarios such as biomedical applications. Concept Bottleneck Models (CBMs) (Kumar et al., 2009; Lampert et al., 2009; Koh et al., 2020) stand out in the landscape of interpretable models as they learn a set of high-level intermediate concepts associated with an outcome, in a manner which facilitates human oversight (Banerji et al., 2025). CBMs first predict the level of activation of each concept for a given input data point, then use these activations to predict a task. Their structure makes them self-explainable, and allows for direct human supervision and intervention, as opposed to post hoc interpretability approaches.
@@
A trained concept-based model, however, may be significantly less interpretable than expected, despite presenting the illusion of being so. A key reason for this is the phenomenon of information leakage (Kazhdan et al., 2021; Margeloiu et al., 2021; Mahinpei et al., 2021; Havasi et al., 2022; Makonnen et al., 2025). Leakage occurs when a poorly designed concept-based model uses additional information to attain a higher task accuracy than a well-designed (i.e. transparent) concept-based model, at the cost of undermining concept interpretability. Storing such additional information in the learnt concepts is beneficial for the model with respect to loss function minimisation, as it typically allows the final head to leverage this information, attaining higher task performance while preserving high accuracy in concept predictions. This ultimately results in learnt concepts not being aligned with the ground-truth concepts, thus hindering the actual interpretability of the model. For example, in the clinical setting of cancer diagnosis, the grade of a tumour (how aggressive it looks under a microscope) is correlated to tumour stage (how much the cancer has spread through the body) and leakage may occur between learned concept representations of grade and stage. This means that a concept representation of grade might also contain information about tumour stage. However, these concepts have independent clinical definitions and their appropriate assessment is essential to guide clinical management. A concept-based model which displays leakage may therefore be unsafe for critical decision-making applications, such as clinical scenarios.
@@
We then employ these measures to identify and assess the primary causes of leakage in CBMs, including: over-expressive concept representations; an incomplete set of annotated concepts; a misspecified final head; and insufficient concept supervision. Building on this novel information-theoretic framework, we provide strong evidence that CEMs and closely related models exhibit dramatic leakage regardless of the choice of hyperparameters, thereby casting serious doubt on their interpretability. Finally, we propose a set of guidelines for the design of concept-based models based on preventing leakage, and we argue that evaluating leakage is a crucial step of model development to ensure interpretability.
The complete codebase for computing leakage scores, evaluating concept-based models, and reproducing all experimental results is publicly available at https://github.com/enricoparisini/xai-concept-leakage.
@@ 2 Concept bottleneck models
Consider a dataset $\left\{x^{(n)},c^{(n)},y^{(n)}\right\}_{n=1}^{N}$ with $N$ observations, where $x^{(n)}\in\mathbb{R}^{m}$ is the input, $c^{(n)}\in\{0,1\}^{k}$ represents a set of $k$ annotated concepts, and $y^{(n)}\in\{1,\dots,\ell\}$ is the target label. A CBM (Figure 1) is the composition of a concept encoder $g^{*}$ mapping the input to the predicted concept activations $\hat{c}=g^{*}(x)$, and a classifier head $f^{*}$ which maps the concepts to the predicted label $\hat{y}=f^{*}(\hat{c})$. This framework easily extends to regression tasks $y$ and to categorical or continuous concepts.
The presence of a concept bottleneck displaying the predicted concept activations for each input allows a human overseer to directly supervise the model reasoning without resorting to post-hoc approaches to explainability such as CAM (Zhou et al., 2016), GradCAM (Selvaraju et al., 2017), LIME (Ribeiro et al., 2016) or SHAP scores (Lundberg and Lee, 2017). Additionally, this setup enables human interventions: if one or more predicted concepts are identified as incorrect for a given input $x^{(n)}$, a human overseer can update their values according to their expertise, resulting in an improved task prediction.
To achieve full interpretability, the final head architecture should represent a functional dependence that is interpretable. This allows for a higher level of supervision, e.g. in terms of which concepts are more relevant for the task prediction. For this reason, linear models are popular.
@@
  • independent training, where the concept encoder and the final head are trained independently on ground-truth data. In this case,
@@
  • sequential training, where the concept encoder is trained on ground-truth data, while the final head is trained on the predicted concepts. In this case,
@@
  • joint training, where the model is trained end-to-end. In this case,
@@
    where $\lambda\geq 0$ is a hyperparameter that quantifies the relative importance of concept and task learning. $\lambda=0$ corresponds to a black box model.
@@
Common choices for concept encoding include binary probabilities, soft probabilities or logits. We will refer to hard CBMs as independently trained models with binary concept encoding, while soft CBMs denote jointly trained models with soft probability-based concept activations. Additionally, we will refer to jointly trained models using logit-based concept encoding as logit CBMs.
@@
More recently, CEMs (Figure 1) were proposed in Espinosa Zarlenga et al. (2022). Here a pair of vectors $\left(\bm{\hat{c}^{+}}_{i},\bm{\hat{c}^{-}}_{i}\right)\in\mathbb{R}^{2d}$ is predicted for each input and concept $c_{i}$, $i=1\dots k$ (where the hyperparameter $d$ indicates the embedding dimension). This pair of vectors is subsequently used to predict the activation $\hat{c}_{i}$ as a soft probability. The weighted vectors $\bm{\hat{c}^{w}}_{i}=\hat{c}_{i}\,\bm{\hat{c}^{+}}_{i}+(1-\hat{c}_{i})\,\bm{\hat{c}^{-}}_{i}$ are then constructed and concatenated before being fed into the final head to predict the label. In this setup $\bm{\hat{c}^{+}}_{i}$ and $\bm{\hat{c}^{-}}_{i}$ are meant to represent concept $i$ being active or inactive respectively. Learning a two-fold representation for each concept enables interventions, by modifying the concept activations $\hat{c}_{i}$ and thus selecting either $\bm{\hat{c}^{+}}_{i}$ or $\bm{\hat{c}^{-}}_{i}$ as an input for the final head.
@@
Training in CEMs is performed end-to-end. Similarly to joint CBMs, the loss function consists of reconstruction losses for both the concept activations and the class label, weighted by a parameter $\lambda\geq 0$ as in (3). To make them more receptive to interventions (i.e. to realize any improvements in task performance upon intervention), CEMs are exposed to interventions during training: each activation $\hat{c}_{i}$ is set to its ground-truth value $c_{i}$ with probability defined by a parameter $p_{int}\in(0,1)$.
@@
Reasoning in CEMs takes place at the level of the vectors $\bm{\hat{c}^{w}}_{i}$, as opposed to CBMs where it is based on concept activations $\hat{c}_{i}$. However, as high-dimensional objects trained only to be predictive of concepts $c_{i}\in\{0,1\}$, the vectors $\bm{\hat{c}^{w}}_{i}\in\mathbb{R}^{d}$ are excessively flexible representations, capable of encoding an arbitrary amount of information about the task label. We demonstrate in Section 8 that this is the fundamental reason why CEMs generally fail to be interpretable for any choice of their hyperparameters. While Espinosa Zarlenga et al. (2022) present CEMs as interpretable models that resolve the accuracy-interpretability trade-off, we argue that they are still affected by this trade-off: indeed, they favour accuracy over interpretability. Although requiring a case-by-case evaluation, other CEM-based architectures, such as Espinosa Zarlenga et al. (2023b); Ridzuan et al. (2024); Gao et al. (2024); Hu et al. (2024); Srivastava et al. (2024), are likely to face similar interpretability challenges.
@@ 3 Definition of leakage
Figure 2: Distributions of predicted activations for concept 1 on the test set in CBMs with increasing leakage from left to right. Colours are based on the ground-truth values of concepts and task label, and to improve visibility, bars with lower counts are rendered in front of those with higher counts.
@@
The explainability of CBMs and CEMs assumes that learnt concepts are aligned with ground-truth human-interpretable concepts. However, on general grounds that is not the case. During training, concept-based models often encode additional input information into the learned concepts to enhance task accuracy. When this happens information leakage occurs (Kazhdan et al., 2021; Margeloiu et al., 2021; Mahinpei et al., 2021; Havasi et al., 2022; Lockhart et al., 2022; Ragkousis and Parbhoo, 2024), and the model does not rely solely on the value of concepts for its task prediction, but also on leaked information, which hinders interpretability. The reasoning of the model thus becomes effectively obscure to human overseers, to a degree specified by the amount of leakage.
@@
In practice, information leakage manifests as additional structure arising in the learnt concept distributions, as shown in Figure 2. In this example models must learn two binary concepts for a binary classification task (see Appendix A.1 for more details). We compare the distributions of activations for the first concept by a hard CBM and two soft CBMs, with either intermediate ($\lambda=1$) or low ($\lambda=0.01$) concept supervision. By definition no leakage can be present in hard CBMs, since concept learning is independent of task learning, and concepts are converted to binary values before being passed to the final head, preventing leaked information. Soft CBMs with lower concept supervision are instead more prone to leakage as accurate concept learning is not enforced during training (see Koh et al., 2020; Kazhdan et al., 2021, and our results in Section 7).
@@
Neglecting the peak close to 0.5, which corresponds to observations the model is most uncertain about, the concept distribution in the hard CBM is very close to the binary ground-truth distribution for concept 1. In the soft CBM with $\lambda=1$ additional structure is learnt based on the value of the other concept and of the class label, while for $\lambda=0.01$ the concept itself is a very strong predictor of the value of both the class label and the other concept. In particular, the distribution is more similar to the ground-truth binary task distribution than to the concept distribution.
@@
Based on these observations, we identify two main types of leakage that concept-based models suffer from, and propose the following classification:
@@
1) Concepts-task leakage: additional information about the task is stored into the learnt concepts. For training to be successful, the annotated concepts must be predictive of the task label as measured by the ground-truth mutual information (MI, see more below) between each concept and the task label. If concepts-task leakage is present, then the learned concepts-task MI will be higher than ground-truth concepts-task MI. This is the predominant type of leakage and the main cause of non-interpretability, as it directly allows the model to achieve a better task performance (see experiments in Section 6).
@@
2) Interconcept leakage: additional information about the other concepts is stored into each concept. In a given dataset, concepts are predictive of the value of other concepts as determined by their ground-truth pair-wise MI. This type of leakage manifests as learned concepts being more predictive of the value of the other learned concepts, resulting in a learned interconcept MI being higher than the ground-truth. This effect is usually a secondary factor of non-interpretability (see experiments in Section 6), but it may be beneficial for models as an internal error-correcting tool: interconcept leakage provides redundant pathways for the model to attain high task performance even when concept predictions are poor (Heidemann et al. (2023)).
@@ 4 Existing measures of interpretability and their shortcomings
Previous works (Koh et al., 2020; Kazhdan et al., 2021; Margeloiu et al., 2021; Mahinpei et al., 2021; Havasi et al., 2022; Espinosa Zarlenga et al., 2023a) have proposed several measures of leakage, however none take into account the information-theoretic nature of leakage outlined in Section 3. As we highlight in this section, these existing measures consequently exhibit inherent limitations in sensitivity and robustness, limiting their usefulness in most practical scenarios.
@@
Although performance metrics such as concept accuracy, F1 score and AUC are general quality indicators of concept learning, they are not sensitive to more subtle effects undermining interpretability, such as leakage. In particular, two models may exhibit the same evaluation scores while considerably differing in terms of leakage. This is apparent from the example in Figure 2, where a classification threshold of 0.5 yields essentially the same concept accuracy and F1 score for the hard and $\lambda=1$ soft CBM, while the latter encodes a non-trivial amount of additional structure. Concept AUC is a finer measure of interpretability than accuracy and F1 score as it captures more information about the concept distribution, and in this simple example it is able to discriminate between those two models. However it is not generally sensitive enough to quantify leakage in more complex setups as we illustrate in Section 6.
@@
These scores were defined in Espinosa Zarlenga et al. (2023a). The Oracle Impurity Score (OIS) is obtained by training $k(k-1)$ additional neural networks (NNs) $\psi_{i,j}$ (typically 2-layer perceptrons of modest size) predicting the value of the ground-truth concept $j$ from the learnt representation of concept $i$. The $k\times k$ impurity matrix defined as ${\pi_{ij}(\hat{c},c)=\textrm{AUC}\left(\psi_{i,j}(\hat{c}_{i}),c_{j}\right)}$ is meant to estimate the predictivity of each learnt concept representation with respect to any other across the dataset. The OIS is the average difference between the predictivities of learnt and ground-truth concepts over pairs of concepts,
@@
where $||\cdot||_{F}$ indicates the Frobenius norm, and the normalisation ensures $0\leq\text{OIS}\leq 1$. The OIS is thus intended to provide an estimate of interconcept leakage – although, it is unable to account for concepts-task leakage, which is the prevalent effect as we illustrate in Section 6. Although demonstrated to be superior to other concept-quality metrics in Espinosa Zarlenga et al. (2023a), the OIS is subject to additional limitations that hinder its applicability. In particular, in Section 6 we demonstrate that:
@@
    its value is typically non-vanishing and relatively high for hard CBMs, indicating bias and a lack of sensitivity;
@@
Figure 3: Performance upon random intervention of the three models analysed in Figure 2. See Appendix A.3 for more details.
The Niche Impurity Score (NIS) was defined to capture leakage manifesting as additional information about concept $j$ stored into a joint subset of the learnt representations $\{i_{1},\dots,i_{p}\}$ not containing concept $j$ (see Espinosa Zarlenga et al., 2023a, for details). This score is meant to go beyond pair-wise interconcept leakage and estimate (generally subordinate) higher-order effects. As illustrated in Appendix B, the NIS is generally non-vanishing and very high for hard CBMs, and moreover it appears to be anticorrelated with intervention performance and leakage, casting doubt on its unsuitability as a measure of leakage.
@@
Interventions can be used to evaluate the interpretability of a model. They are performed by substituting the ground-truth values of concepts to the corresponding predicted activation, with a given policy defining the order of the concepts to intervene on. The behaviour of the task accuracy after each intervention measures the extent to which the learnt and ground-truth concepts are aligned. A task accuracy that decreases is a coarse indication of leakage since it suggests that the final head expects additional information that is not present in the ground-truth concept distribution. Figure 3 shows the intervention performance of the same three models as in Figure 2. As expected, models with higher leakage have worse intervention performance, while the hard CBM displays a monotonically increasing task accuracy.
As a metric of leakage, intervention performance has several limitations, however:
@@
    Because it is based on evaluation scores, it does not address the information-theoretical nature of leakage, resulting in a limited resolution. It is unable to distinguish between interconcept and concepts-task leakage, and it does not directly quantify interpretability at the level of individual concepts. Typically it is therefore not sufficiently informative to guide model design or hyperparameter selection.
@@
    Its evaluation is computationally expensive, especially in real-world applications in which models have a large number of concepts and are evaluated using state-of-the-art policies with a high overhead such as CooP (Chauhan et al., 2023).
@@
    In more complex architectures such as CEMs, it is not a measure of interpretability, see Appendix G.
Collectively these issues indicate that poor intervention performance is a measure of leakage with perfect specificity (leakage has always occurred when intervention performance is poor), but with poor sensitivity (when intervention performance is good, leakage is not guaranteed to be absent). Although the leakage scores we propose in Section 5 address these issues, intervention performance generally still remains a coarse but useful indicator of leakage in CBMs.
Let us denote by $y_{acc}^{*(k)}$ the task accuracy after all the $k$ concepts have been intervened on in the model under evaluation, which coincides with the accuracy of the final head on the ground-truth concepts. As a benchmark of the proposed leakage scores we will use the intervention score
$S_{int}=y_{acc}^{(k)}-y_{acc}^{*(k)}\,,$ (5)
where $y_{acc}^{(k)}$ is the reference accuracy of the final head when separately trained on the ground-truth concepts. It represents the maximal task accuracy attainable by the final head and it measures the fundamental misspecification of the final head with respect to the ground-truth functional dependence $y=f(c)$. $S_{int}$ thus quantifies the further decrease in task accuracy caused by leakage accounting for final head misspecification. Its value is independent of the intervention policy adopted. Note that $y_{acc}^{(k)}$ coincides with $y_{acc}^{*(k)}$ for a successfully trained hard model, thus $S_{int}=0$ for hard models by construction.